OpenStudy (anonymous):

3(a + b) = 3a + 3b is an example of which algebraic property? Distributive Property Associative Property of Addition Commutative Property of Addition Symmetric Property

OpenStudy (anonymous):

wow for a change it is not C go with Distributive Property

OpenStudy (anonymous):

you can tell because it has both \(+\) and \(\times \) in it in english (well math english) you say "multiplication distributes over addition" i.e \(a(b+c)=ab+ac\)